# PhyT2V: LLM-Guided Iterative Self-Refinement for Physics-Grounded Text-to-Video Generation

### v1

@@ B.2 Details of Prompt Refinement in PhyT2V
In this subsection, we provide more detailed examples of the process of prompt refinement in PhyT2V, with the CogVideoX-5B T2V model on the VideoPhy and PhyGenBench prompt datasets. Fig 19, 20, and 21 show the generated videos in categories of solid-to-solid, solid-to-fluid, and fluid-to-fluid for the VideoPhy dataset. Fig 22, 23 and 24 show the generated videos in categories of mechanics, optics and thermal on the PhyGenBench dataset.
Appendix C Details of the Ablation Study
C.1 The Impact of T2V Model size
@@
By comparing the videos generated by the CogVideoX-2B and CogVideoX-5B models with the same text prompt, we found that the PhyT2V approach can unleash more power of physical-grounded video generation and achieve better quality of the generated video, with a T2V model with larger parameter sizes. Results are show in Figure 25.
C.2 Components in Prompts
In this section, we remove some components from the prompts being used in two parallel subproblems and the final subproblem described in Section 3.1 and 3.2 of the main texts of the paper, and investigate how such removal affects the video generation. Results are shown in Figure 26, 27 and 28, respectively. These results show that, without the first sentence of role indicator (“You are a physics expert.”), the generated output content is lake of precise information, On the other hand, when the in-context examples ($[E]$) are missing, the LLM can not generate the output in the expected format.
Appendix D Difficult T2V Cases
Finally, we also found that, even when using the PhyT2V to refine the T2V process for multiple rounds, the generated video still failed to correctly reflect some real-world common knowledge. For example, Figure 29 shows that the generated video would fail to precisely generate the human body, especially the human hands. We believe that PhyT2V’s difficulty in these cases is consistent with the limitations of the existing T2V models, and can only be addressed by further improving the design of T2V model architectures.
Figure 11: Examples of generated videos in the category of fluid to fluid in the VideoPhy dataset
Figure 12: Examples of generated videos in the category of solid to fluid in the VideoPhy dataset
Figure 13: More examples of generated videos in the category of solid to fluid in the VideoPhy dataset
Figure 14: Examples of generated videos in the category of solid to solid in the VideoPhy dataset
Figure 15: More examples of generated videos in the category of solid to solid in the VideoPhy dataset
Figure 16: Examples of generated videos in the category of mechanics in the PhyGenBench dataset
Figure 17: Examples of generated videos in the category of optics in the PhyGenBench dataset
Figure 18: Examples of generated videos in the category of thermal in the PhyGenBench dataset
Figure 19: An example showing details of PhyT2V’s prompt refinement, in the category of solid-to-solid in the VideoPhy dataset
Figure 20: An example showing details of PhyT2V’s prompt refinement, in the category of solid-to-fluid in the VideoPhy dataset
@@ Appendix D Difficult T2V Cases
Figure 21: An example showing details of PhyT2V’s prompt refinement, in the category of fluid-to-fluid in the VideoPhy dataset
@@
Figure 22: An example showing details of PhyT2V’s prompt refinement, in the category of mechanics in the PhyGenBench dataset
@@
Figure 23: An example showing details of PhyT2V’s prompt refinement, in the category of optics in the PhyGenBench dataset
@@
Figure 24: An example showing details of PhyT2V’s prompt refinement, in the category of thermal in the PhyGenBench dataset
Figure 25: The impact of T2V model size on the quality of generated video
@@
Figure 27: The prompt without the sentence of role indicator (“You are a physics expert.”) and output
Figure 28: The prompt without the in-context examples and output
Figure 29: Examples of difficult T2V cases of generating human hands

### v2

@@ B.2 Refinement process details
In this subsection, we present the prompt refinement detail by the CogVideo-5b with PhyT2V on the VideoPhy and PhyGenBench dataset, Fig 19, 18, 17 show the prompt refinement details of fluid to fluid, solid to fluid, solid to solid cases on VideoPhy dataset, Fig 20, 21, 22 show prompt refinement details of force, optics and thermal cases on PhyGenBench dataset.
Appendix C Ablation study details
C.1 Model size
@@
We found that the PhyT2V approach can unleashing more power of physical-grounded video generation on a larger model as the result shown by comparing the CogVideo-2b and CogVideo-5b in Figure 23 .
C.2 Prompt template component
In this section some part of the prompt template component is removed to show the necessity of the corresponding components as shown in Figure24, 25, 26. Without the role indicator sentence, the generated output content is lake of precise information, without the in-context examples, the GPT can not generated the output in an expected format.
Appendix D Failure cases
PhyT2V may be ineffective in two categories of cases.
First, many T2V models exhibit temporal flickering or inconsistent object trajectories, due to absence of long-term temporal coherence mechanisms in model design. Even if prompts are refined to emphasize smooth temporal transitions or continuous motions, these requirements may not be achieved due to model’s limitations.
Second, T2V models are typically trained on large datasets, which often lack samples of rare or complex physical phenomena. Hence, these models struggle in scenarios that are underrepresented in the training data. Even with highly specific prompts, the T2V models may still fail to extrapolate effectively to these underrepresented cases.
For some specific generation content categories, we found that even with the PhyT2V refined several rounds, the diffusion model still failed to precisely generating human body, especially on human hands as shown in Figure 27.
@@ Appendix D Failure cases
Figure 11: Video generation example on fluid to fluid specific prompt in VideoPhy dataset
Figure 12: Video generation example on solid to fluid specific prompt in VideoPhy dataset
Figure 13: Video generation example on solid to solid specific prompt in VideoPhy dataset
Figure 14: Video generation example on force specific prompt in PhyGenBench dataset
Figure 15: Video generation example on optics specific prompt in PhyGenBench
Figure 16: Video generation example on thermal specific prompt in PhyGenBench
Figure 17: Refinement detail example on solid to solid specific prompt in VideoPhy dataset
Figure 18: Refinement detail example on solid to fluid specific prompt in VideoPhy dataset
Figure 19: Refinement detail example on fluid to fluid specific prompt in VideoPhy dataset
@@
Figure 20: Refinement detail example on force specific prompt in PhyGenBench dataset
@@
Figure 21: Refinement detail example on optics specific prompt in PhyGenBench dataset
@@
Figure 22: Refinement detail example on thermal specific prompt in PhyGenBench dataset
Figure 23: Ablation study on model size
@@
Figure 25: The prompt without the role indicator and output
Figure 26: The prompt without the in-context example and output
Figure 27: Failure case example on human hand related video generation